# Stronger Random Baselines for In-Context Learning

### v1

@@ 1 Introduction
Downstream users want to find and deploy the best prompt (Mizrahi et al., 2023). But ICL performance of LMs varies greatly across semantically equivalent prompt features like the choice of demonstrations and their order (Zhao et al., 2021; Lu et al., 2022), instruction phrasing (Mizrahi et al., 2023), and template formatting (Sclar et al., 2024; Voronov et al., 2024). Because performance is both variable and unpredictable, standard ICL practice involves searching over large numbers of potential prompts based on a validation set. Researchers often report model performance as the maximum score on a validation set or the corresponding performance of that prompt on an additional truly held-out test set to avoid overfitting to the validation set (Brown et al., 2020; Perez et al., 2021). In this work we show that we can better identify prompts that may be overfitting and avoid unneccessary evaluations on test data.
@@
In searching for the best prompts, the simplest and most common comparison is to a random baseline. The standard random baseline for classification tasks is the expected accuracy of guessing labels uniformly at random (Mitchell, 1997, inter alia). While universally treated as a point estimate, the accuracy of any specific random classifier follows a binomial distribution. For larger datasets, the variance of this distribution is tightly concentrated around the expectation, but variance can be considerably higher for the small datasets typically used in the ICL setting. We introduce a stronger random baseline that accounts for both variance and validation set reuse by asking a fairer question: if we are choosing the best of different prompts, why not compare that prompt’s accuracy to the best of different random classifiers? Figure 1 shows how these two random baselines can lead to different conclusions.
@@
Treating random performance as a distribution has two key advantages. First, the stronger random baseline can be calculated in closed form as the expectation of the maximum order statistic of the binomial distribution. When choosing the best prompt from even as few as 10 options, this baseline increases the threshold for beating “random” performance by more than 7 points of accuracy for a binary classification task with 100 examples. Second, by using a family of parametric distributions to represent random performance rather than a point estimate, we can also calculate a comparable performance metric across multiple datasets that takes into account factors like the number of classification options, the number of evaluated prompts, and the number of evaluated examples. We can then use the tail probability of the random distribution as a -value to quantify what fraction of random classifiers outperform a prompt. This contextualization also applies in the true few-shot setting, where evaluation data is used only once.
@@
Our contributions are the following. First, we propose a maximum random baseline that explicitly depends on validation set size and the number of evaluated prompts. We give a simple method for computing this value by taking the expectation of the maximum order statistic of the binomial distribution and show how this quantity varies with task parameters and evaluation setup. Second, we examine quantized LM few-shot performance across the BIG-bench Lite benchmark suite when choosing the best prompt demonstrations, finding in our setting that more than 20% of results that exceed the standard baseline do not exceed the maximum random baseline. Third, we show that when a truly held-out test set is available, comparing maximum validation performance to the maximum random baseline, rather than the standard random baseline, is better able to predict whether corresponding test performance will also be above random. This can help prevent prematurely evaluating with a held-out set, reducing the chance of overfitting to that held-out test set. The stronger baseline can also allow researchers to use smaller validation sets more confidently. Finally, the maximum random baseline can be easily calculated as a drop-in replacement for the standard random baseline.111Code is available at: https://github.com/gyauney/max-random-baseline High-quality datasets remain rare and expensive. They will be reused, and our evaluation baselines should account for this.
@@ 2 Related work
Data reuse has been approached from many angles. Some restrict access to the test set, either through submission systems (Wang et al., 2019a; b; Alex et al., 2021; Bragg et al., 2021) or differential privacy (Dwork et al., 2015). Perez et al. (2021) advocate for a true few-shot setting: model selection using cross-validation and minimum description length on the training set, with only one held-out set evaluation. While they compare selected prompts to randomly selected prompts, we instead compare to classifiers that guess at random. Some works create held-out sets for ICL (Wei et al., 2022; Nori et al., 2023), though this is not always standard practice. Rather than stipulate how to access the validation set, our approach contextualizes the current practice of reusing the validation set. Sometimes the prompt template with the best validation performance is chosen for test set evaluation (Wei et al., 2022; Mizrahi et al., 2023). The stronger random baseline does not apply to the test accuracy, but it does apply to the best validation accuracy.
@@ 3 Random baselines
Consider a dataset with validation examples and possible labels.222Appendix A extends the baseline to datasets where varies per example. Then is the probability of guessing one example’s label correctly uniformly at random. We primarily consider how to contextualize accuracy in a setting with just one set of examples (generalization is discussed in Section 5.2). An experiment evalutes different classifiers (or prompts, or hyperparameter settings) and reports maximum accuracy.
@@
-values against random baselines.
For either kind of random baseline, it can be useful to compare to the distribution of random performances rather than just the expected accuracy. -values are a general tool for calculating the probability that a value drawn from a distribution is the same as or more extreme than a given value (Wasserman, 2013).333Note that we use this term divorced from any hypothesis testing framework. The -values for the accuracy of a classifier against both baselines are (details in Appendix A):
@@ 4 Properties of the expected maximum random baseline
When (the true few-shot setting), the expected maximum is simply the expectation of the binomial distribution. The maximum random baseline therefore subsumes the standard baseline as a special case. When is large, the expected maximum random accuracy is nearly the same as the standard random baseline because the binomial distribution is very concentrated around its expectation with large . The implication of this result is that if an experiment has a large , the standard random baseline is only a good indicator of random guessing if is also large. But while increasing is always a good idea theoretically, in practice it may not be feasible for the kind of diverse, fast-moving LM tasks that are the key use case for ICL.
@@ 5.1 The standard random baseline overstates performance
Figure 3 shows OLMo-7B performance for selected tasks. The -axis shows the number of prompts evaluated on the validation set. As increases, the maximum random baseline also increases, sharply at first and then more gradually. For emoji_movie and known_unknowns, the expected maximum 1-, 2-, and 4-shot accuracies across multiple prompts are above the standard random baseline but below the maximum random baseline. To further illustrate how the baselines disagree, we can use their distributions to generate -values for the observed accuracies. In the middle row of panels, -values with respect to the standard random distribution are near-zero in almost all cases. The bottom row shows -values with respect to the maximum baseline: here 1-, 2-, and 4-shot all have high -values. Contextualizing performance against the maximum random baseline shows that these datasets are more challenging than they appear. Performance is high enough on tasks like bbq_lite_json and hindu_knowledge to be above random no matter which baseline is used. In these cases, -values for both baselines are near-zero. The maximum random baseline gives us additional confidence that performance is in fact good.
Figure 4 compares base and instruction-tuned models on the emoji_movie dataset. These results show that the standard random baseline substantially underestimates the difficulty of this dataset. Comparing instead to the maximum random baseline reveals that only Llama-2-7b outperforms random guessing. While the instruction-tuned falcon-7b models slightly outperform the standard random baseline, the standard -values show that a significant percentage of random classifiers still have higher accuracy.
@@
Table 1 shows that out of 288 total experiments, maximum validation accuracy exceeded the standard random baseline in 255, but maximum validation accuracy exceeded the maximum random baseline in 199. The baselines disagreed in 56 experiments. This means that of results that are above the standard baseline are not above the maximum baseline. Among base models, there are fewer baseline disagreements as the number of shots increases. Surprisingly, there are not fewer disagreements when moving from base models to their instruction-tuned counterparts. Specific datasets are responsible for many disagreements, the highest number of which come from formal_fallacies and language_identification. For many datasets, both baselines agree for all evaluations. Full per-dataset results are in Figures 11, 12, and 13 in Appendix D.
@@ 5.2 Maximum random baseline predicts held-out accuracy
In the previous section we focused on contextualizing maximum validation accuracy, which is commonly reported for ICL tasks (Perez et al., 2021). Validation accuracy is often used as a first step to select a model, followed by reporting the selected model’s performance on an additional held-out test set. But the more we look at the test set, the less useful it becomes. Here we show that the maximum random baseline is a better predictor of whether test performance will exceed random guessing than the standard random baseline and thus can avoid wasted test set evaluations.
@@
Just as before, we choose the prompt with best performance on the validation data and report: the maximum validation accuracy, the standard random baseline, and the maximum random baseline. Given just the validation accuracy, we have to make a decision: is this prompt’s test accuracy above or below random? Using the same models and 16 BIG-bench Lite tasks as above, we randomly partition each task into a validation set (75%) and a test set (25%). For a given model and task, we select the prompt from among with the highest validation accuracy. This prompt’s validation accuracy is compared to the standard random and the maximum random baselines.
The prompt is said to outperform random on the corresponding held-out test set when its test accuracy is above the standard random baseline (because there has been no test set reuse). For each of 100 different splits of each dataset, we use validation accuracy to predict whether test set accuracy is above standard random. We are comparing two “classifiers”: standard predicts true when validation accuracy is above standard random, and max predicts true when validation accuracy is above the maximum random baseline. In both cases it is possible that a low quality prompt could have high validation accuracy due to random chance, so we do not expect perfect performance. Another reason we cannot expect perfect performance on this task is that just knowing that validation performance is above the standard random baseline means that performance is better than that of at least half of the random classifiers, not that it outperforms all of them. But we can evaluate which baseline gives us more insight into held-out accuracy.
@@ Results.
Evaluating all predictions across all splits of datasets, the held-out accuracy of the best prompt exceeds random 73% of the time. Figure 5 shows that when predicting whether test performance is above random from validation accuracy, standard achieves an accuracy of 0.82, , and . max has a lower accuracy of 0.79, but does better on and . standard achieves higher accuracy on this task at the expense of many false positives—it often incorrectly predicts that test performance will be above random chance. In fact, the standard random baseline does only slightly better in this setting than simply predicting that all test accuracies will be above random. The maximum random baseline has higher precision—when validation accuracy surpasses the maximum random baseline we can be more confident that the corresponding held-out performance is above random. Results split by model and dataset are in Appendix C, as are different values of for validation set reuse. As the number of prompts that are being searched over increases, the maximum random baseline maintains a lower false positive rate even as the standard baseline’s false positive rate increases (Figure 10, Appendix C).
We can additionally use each baseline’s distribution function as a measure of how likely validation performance is to be above random. Instead of directly predicting above or below random as compared to each baseline’s expected accuracy, we instead associate each point with the percentage of random classifiers that the validation accuracy is above. Both baselines produce the same rankings of likelihood above the baseline. This is because the distribution function of the maximum order statistic of the binomial distribution is a monotonic function of the distribution function of the binomial distribution (Equations 3, 6). Both baselines therefore yield the same ROC and precision-recall curves (, ). The threshold for predicting above or below random performance is what changes, leading to the difference shown in Figure 5. In both cases, using the additional information provided by treating the random baselines as distributions contextualizes performance better than point estimates alone.
@@ 6 Discussion and conclusion
If an experiment is calculating the maximum performance across multiple evaluations on a dataset, the standard random baseline is not necessarily a good representation of the probability of achieving that performance by random guessing. ICL is particularly susceptible to this danger because of its combination of small validation sets, many prompt evaluations, and difficult tasks. In such settings, the expected maximum accuracy across multiple random classifiers is a stronger and more appropriate baseline. This metric is easily calculated, and we release code for a drop-in replacement baseline.
While it is best practice to report test performance on a truly held-out set of examples (Perez et al., 2021), the maximum random baseline offers additional contextualization of maximum validation accuracy. It is common to report validation size , and reporting the number of prompt evaluations needed for this baseline is also good practice. Of a random sample of 20 papers from EMNLP 2023 that study ICL, six report using multiple prompts, three state that they report the maximum evaluation accuracy over multiple prompts, and only one of these reports how many times the validation set was used. We argue that reporting such settings allows for proper contextualization of results and should be a best practice.
We reiterate the many calls for larger evaluation sets (Card et al., 2020; Bragg et al., 2021), this time to limit the variance of random guessing. But the reality of contemporary LM evaluation is that researchers have limited time and budget for computation and dataset development. Producing thousands of high-quality examples and evaluating many possible prompts on them may not always be feasible (Liang et al., 2023; Polo et al., 2024). The maximum random baseline proposed in this paper enables researchers to reduce validation size and therefore computation, while still avoiding falsely positive prompt settings.
We leave it to future work to study the maximum random baseline across more models, tasks, prompt variations, and scoring strategies, as well as extensions to metrics beyond accuracy like F1. Future work can also study the rigorous specification of hypothesis tests for deciding whether generalization performance is likely to exceed random guessing. While we limit our analysis to in-context learning with language models, the stronger random baseline applies to any classification setting with evaluation set reuse.

### v2

@@ 1 Introduction
Downstream users want to find and deploy the best prompt (Mizrahi et al., 2024). But ICL performance of LMs varies greatly across semantically equivalent prompt features like the choice of demonstrations and their order (Zhao et al., 2021; Lu et al., 2022), instruction phrasing (Mizrahi et al., 2024), and template formatting (Sclar et al., 2024; Voronov et al., 2024). Because performance is both variable and unpredictable, standard ICL practice involves searching over large numbers of potential prompts based on a validation set. Researchers often report model performance as the maximum score on a validation set or the corresponding performance of that prompt on an additional truly held-out test set to avoid overfitting to the validation set (Brown et al., 2020; Perez et al., 2021). In this work we show that we can better identify prompts that may be overfitting and avoid unneccessary evaluations on test data.
@@
In searching for the best prompts, the simplest and most common comparison is to a random baseline. The standard random baseline for classification tasks is the expected accuracy of guessing labels uniformly at random (Mitchell, 1997, inter alia). While universally treated as a point estimate, the accuracy of any specific random classifier follows a binomial distribution. For larger datasets, the variance of this distribution is tightly concentrated around the expectation, but variance can be considerably higher for the small datasets typically used in the ICL setting. We introduce a stronger random baseline that accounts for both variance and validation set reuse by asking a fairer question: if we are choosing the best of different prompts, why not compare that prompt’s accuracy to the best of different random classifiers? Figure 1 shows how these two random baselines can lead to different conclusions on a fixed evaluation set. We also find that the maximum baseline can provide a more accurate estimate of a prompt’s generalization to a held-out test set than the standard baseline does. This can reduce premature usage of the test set.
@@
Treating random performance as a distribution has two key advantages. First, the stronger random baseline can be calculated in closed form as the expectation of the maximum order statistic of the binomial distribution. When choosing the best prompt from even as few as 10 options, this baseline increases the threshold for beating “random” performance by more than 7 points of accuracy for a binary classification task with 100 examples. Second, by using a family of parametric distributions to represent random performance rather than a point estimate, we can also calculate a comparable performance metric across multiple datasets that takes into account factors like the number of classification options, the number of evaluated prompts, and the number of evaluated examples. We can then use a tail probability to quantify what fraction of random classifiers outperform a prompt. This contextualization also applies in the setting where evaluation data is used only once.
@@
Our contributions are the following. First, we propose a maximum random baseline that explicitly depends on validation set size and the number of evaluated prompts. We give a simple method for computing this value by taking the expectation of the maximum order statistic of the binomial distribution and show how this quantity varies with task parameters and evaluation setup. Second, we examine quantized LM few-shot performance across the BIG-bench Lite benchmark suite when choosing the best prompt demonstrations, finding in our setting that more than 20% of results that exceed the standard baseline do not exceed the maximum random baseline. Third, we show that when a truly held-out test set is available, comparing maximum validation performance to the maximum random baseline, rather than the standard random baseline, is better able to predict whether corresponding test performance will also be above random. This can help prevent prematurely evaluating with a held-out set, reducing the chance of overfitting to that held-out test set. The stronger baseline can also allow researchers to use smaller validation sets more confidently. Finally, the maximum random baseline can be easily calculated as a drop-in replacement for the standard random baseline.111Code is available at: https://github.com/gyauney/max-random-baseline High-quality evaluation datasets remain small, and prompt variability leads to dataset reuse. Our baselines should account for this.
@@ 2 Related work
Data reuse has been approached from many angles. Some restrict access to the test set, either through submission systems (Wang et al., 2019a; b; Alex et al., 2021; Bragg et al., 2021) or differential privacy (Dwork et al., 2015). Perez et al. (2021) advocate for a true few-shot setting: model selection using cross-validation and minimum description length on the training set, with only one held-out set evaluation. While they compare selected prompts to randomly selected prompts, we instead compare to classifiers that guess at random. Some works create held-out sets for ICL (Wei et al., 2022; Nori et al., 2023), though this is not always standard practice. Rather than stipulate how to access the validation set, our approach contextualizes the current practice of reusing the validation set. Sometimes the prompt template with the best validation performance is chosen for test set evaluation (Wei et al., 2022; Mizrahi et al., 2024). The stronger random baseline does not apply to the test accuracy, but it does apply to the best validation accuracy.
@@ 3 Random baselines
Consider a dataset with validation examples and possible labels.222Appendix A extends the baseline to datasets where varies per example. Then is the probability of guessing one example’s label correctly uniformly at random.333Appendix A also shows this does not depend on the dataset having balanced labels. We primarily consider how to contextualize accuracy in a setting with just one set of examples (generalization is discussed in Section 5.2). An experiment evaluates different classifiers (or prompts, or hyperparameter settings) and reports maximum accuracy.
@@
Tail probabilities against random baselines.
For either kind of random baseline, it can be useful to compare to the distribution of random performances rather than just the expected accuracy. To get the probability that random classifiers outperform a given classifier , the tail probabilities for the accuracy of against both baselines are (details in Appendix A):
@@ 4 Properties of the expected maximum random baseline
When (only evaluating one random classifier), the expected maximum is simply the expectation of the binomial distribution. The maximum random baseline therefore subsumes the standard baseline as a special case. When is large, the expected maximum random accuracy is nearly the same as the standard random baseline because the binomial distribution is very concentrated around its expectation with large . The implication of this result is that if an experiment has a large , the standard random baseline is only a good indicator of random guessing if is also large. But while increasing is always a good idea theoretically, in practice it may not be feasible for the kind of diverse, fast-moving LM tasks that are the key use case for ICL.
@@ 5.1 The standard random baseline overstates performance
Figure 3 shows OLMo-7B performance for selected tasks. The -axis shows the number of prompts evaluated on the validation set. As increases, the maximum random baseline also increases, sharply at first and then more gradually. For emoji_movie and known_unknowns, the expected maximum 1-, 2-, and 4-shot accuracies across multiple prompts are above the standard random baseline but below the maximum random baseline. To further illustrate how the baselines disagree, we can use their distributions to generate tail probabilities for the observed accuracies. In the middle row of panels, tail probabilities with respect to the standard random distribution are near-zero in almost all cases. The bottom row shows tail probabilities with respect to the maximum baseline: here 1-, 2-, and 4-shot all have high tail probabilities. Contextualizing performance against the maximum random baseline shows that these datasets are more challenging than they appear. Performance is high enough on tasks like bbq_lite_json and hindu_knowledge to be above random no matter which baseline is used. In these cases, tail probabilities for both baselines are near-zero. The maximum random baseline gives us additional confidence that performance is in fact good.
Figure 4 compares base and instruction-tuned models on the emoji_movie dataset. These results show that the standard random baseline substantially underestimates the difficulty of this dataset. Comparing instead to the maximum random baseline reveals that only Llama-2-7b outperforms random guessing. While the instruction-tuned falcon-7b models slightly outperform the standard random baseline, the standard tail probabilities show that a significant percentage of random classifiers still have higher accuracy.
@@
Table 1 shows that out of 288 total experiments, maximum validation accuracy exceeded the standard random baseline in 255, but maximum validation accuracy exceeded the maximum random baseline in 199. The baselines disagreed in 56 experiments. This means that of results that are above the standard baseline are not above the maximum baseline. Among base models, there are fewer baseline disagreements as the number of shots increases. Surprisingly, there are not fewer disagreements when moving from base models to their instruction-tuned counterparts. Specific datasets are responsible for many disagreements, the highest number of which come from formal_fallacies and language_identification. For many datasets, both baselines agree for all evaluations. Full per-dataset results are in Figures 9, 10, and 11 in Appendix E. Appendix C gives results for non-quantized models.
@@ 5.2 Maximum random baseline predicts held-out accuracy
Now that we have shown that the standard and maximum random baselines differ in many typical ICL settings, we turn to contextualizing held-out test set accuracy. Validation accuracy is often used as a first step to select a model, followed by reporting the selected model’s performance on an additional held-out test set. But, of course, the more we look at the test set, the less useful it becomes. Here we show that the maximum random baseline is a better predictor of whether test performance will exceed random guessing than the standard random baseline and thus can avoid wasted test set evaluations.
@@
Just as before, we choose the prompt with best performance on the validation data and report: the maximum validation accuracy, the standard random baseline, and the maximum random baseline. Given just the validation accuracy, we have to make a decision: is this prompt’s test accuracy above or below random? Using the same models and 16 BIG-bench Lite tasks as above, we randomly partition each task into a validation set (75%) and a test set (25%). For a given model and task, we select the prompt from among prompts with the highest validation accuracy. This prompt’s validation accuracy is compared to the standard random and the maximum random baselines.
The prompt is said to outperform random on the corresponding held-out test set when its test accuracy is above the standard random baseline (because there has been no test set reuse). For each of 100 different splits of each dataset, we use validation accuracy to predict whether test set accuracy is above the standard baseline. We are comparing two “classifiers”: standard predicts true when validation accuracy is above standard random, and max predicts true when validation accuracy is above the maximum random baseline. In both cases it is possible that a low quality prompt could have high validation accuracy due to random chance, so we do not expect perfect performance. Another reason we cannot expect perfect performance on this task is that just knowing that validation performance is above the standard random baseline means that performance is better than that of at least half of the random classifiers, not that it outperforms all of them. But we can evaluate which baseline gives us more insight into held-out accuracy.
@@ Results.
Evaluating all predictions across all splits of datasets, the held-out accuracy of the best prompt exceeds random 73% of the time. Figure 5 shows that when predicting whether test performance is above random from validation accuracy, standard achieves an accuracy of 0.82, , and . max has a lower accuracy of 0.79, but does better on and . standard achieves higher accuracy on this task at the expense of many false positives—it often incorrectly predicts that test performance will be above random chance. In fact, the standard random baseline does only slightly better in this setting than simply predicting that all test accuracies will be above random. The maximum random baseline has higher precision—when validation accuracy surpasses the maximum random baseline we can be more confident that the corresponding held-out performance is above random. Results split by model and dataset are in Appendix D, as are different values of for validation set reuse. As the number of prompts that are being searched over increases, the maximum random baseline maintains a lower false positive rate even as the standard baseline’s false positive rate increases (Figure 8, Appendix D).
We can additionally use each baseline’s distribution function as a measure of how likely validation performance is to be above random. Instead of directly predicting above or below random as compared to each baseline’s expected accuracy, we instead associate each point with the percentage of random classifiers that the validation accuracy is above. Both baselines produce the same rankings of likelihood above the baseline. This is because the distribution function of the maximum order statistic of the binomial distribution is a monotonic function of the distribution function of the binomial distribution (Equations 3, 6). Both baselines therefore yield the same ROC and precision-recall curves (, ). The threshold for predicting above or below random performance is what changes between baselines, leading to the difference shown in Figure 5. In both cases, using the additional information provided by treating the random baselines as distributions contextualizes performance better than point estimates alone.
@@ 6 Discussion and conclusion
If an experiment reports the maximum performance across multiple evaluations on a dataset, the standard random baseline may significantly underestimate the probability of achieving that performance by random guessing. In-context learning is particularly susceptible to this danger because of its combination of small evaluation datasets, many prompt evaluations, and difficult tasks. In such settings, the expected maximum accuracy across multiple random classifiers is a stronger and more appropriate baseline. The baseline is easily calculated, and we release code for a drop-in replacement baseline.
Using the maximum random baseline for calibrating performance on validation sets can avoid inappropriate use of test sets. When a held-out test set is available, the maximum baseline can better predict the generalization performance of the best-performing prompt on the validation set when it is evaluated on a truly held-out test set. This can prevent test-set overuse: if the best prompt doesn’t outperform the maximum random baseline on the validation set, then do not evaluate on the test set yet. The maximum baseline can gracefully account for dataset re-use in other ways, too. Even if an individual researcher strictly uses a benchmark’s test set exactly once, benchmarks are re-used across studies (Koch et al., 2021). The maximum baseline can contextualize results on re-used test sets.
The maximum random baseline highlights the need to report more standard information to contextualize maximum accuracy across prompts. It is common to report validation size , and reporting the number of prompt evaluations needed for this baseline is also good practice. Of a random sample of 20 papers from EMNLP 2023 that study ICL, six report using multiple prompts, three state that they report the maximum evaluation accuracy over multiple prompts, and only one of these reports how many times the validation set was used. We argue that reporting such parameters allows for proper contextualization of results and should be a best practice.
We reiterate the many calls for larger evaluation sets (Card et al., 2020; Bragg et al., 2021), this time to limit the variance of random guessing. But the reality of contemporary LM evaluation is that researchers have limited time and budget for computation and dataset development. Producing thousands of high-quality examples and evaluating many possible prompts on them may not always be feasible (Liang et al., 2023; Polo et al., 2024). The maximum random baseline proposed in this paper enables researchers to reduce validation size and therefore computation and cost, while still avoiding falsely positive prompt settings.
We leave it to future work to study the maximum random baseline across more models, tasks, prompt variations, and scoring strategies, as well as extensions to metrics beyond accuracy, such as F1. Future work can also study the rigorous specification of hypothesis tests for deciding whether generalization performance is likely to exceed random guessing, where dependence between classifiers poses a challenge for analysis. While we limit our analysis to in-context learning with language models, the stronger random baseline applies to any classification setting with evaluation set reuse.